A box contains 10 mangoes out of which 4 are rotten. 2 mangoes are taken out together. If one of them is found to be good, then probability that the other…एक डिब्बे में 10 आम हैं जिनमें से 4 सड़े हुए हैं। 2 आम एक साथ निकाले जाते हैं। यदि उनमें से एक अच्छा पाया जाता है, तो संभावना यह कि दूसरा भी अच्छा है, क्या है?
A box contains 10 mangoes out of which 4 are rotten. 2 mangoes are taken out together. If one of them is found to be good, then probability that the other is also good is
एक डिब्बे में 10 आम हैं जिनमें से 4 सड़े हुए हैं। 2 आम एक साथ निकाले जाते हैं। यदि उनमें से एक अच्छा पाया जाता है, तो संभावना यह कि दूसरा भी अच्छा है, क्या है?
Show answerHide answer
Correct answer: (C) 5/135/13
Explanation
Given at least one good: P(both good) = C(6,2)/C(10,2) = 15/45 and P(at least one good) = 1 − C(4,2)/45 = 39/45. The conditional probability is 15/39 = 5/13.कम से कम एक अच्छा होने पर: P(दोनों अच्छे) = C(6,2)/C(10,2) = 15/45 और P(कम से कम एक अच्छा) = 1 − C(4,2)/45 = 39/45। सप्रतिबंध प्रायिकता 15/39 = 5/13 है।